OpenStudy (anonymous):

find dy/dx by implicit differentiation 3cosxsiny=1

myininaya (myininaya):

product rule is needed for the cos(x)*sin(y) part chain rule is needed for the sin(y) part

myininaya (myininaya):

\[3 \cdot [ \sin(y) \frac{d}{dx}(\cos(x))+\cos(x) \frac{d}{dx}( \sin(y))]=\frac{d}{dx}(1)\]

myininaya (myininaya):

so easy question what is the derivative of (cos(x)) w.r.t x?

OpenStudy (anonymous):

-sinx

myininaya (myininaya):

ok so now we have \[3[\sin(y)(-\sin(x))+\cos(x)\frac{d}{dx}(\sin(y))]=0 \]

myininaya (myininaya):

last part for differentiating what is the derivative of sin(y) w.r.t x

OpenStudy (anonymous):

cosy(y')

myininaya (myininaya):

yep

myininaya (myininaya):

\[3[-\sin(y)\sin(x)+\cos(x)\cos(y)y']=0 \]

myininaya (myininaya):

now you can solve for y' i suggest distributing the 3 first though

OpenStudy (anonymous):

thanks so much these problems seem so scary at first but they are getting easier

myininaya (myininaya):

cool stuff they aren't too scary :)